OpenStudy (bananas1234):

Divide. Assume that no denominator equals zero. 10d^2/4c divided by cd^3/14

OpenStudy (unklerhaukus):

\[\frac{10d^2}{4c}\div\frac{cd^3}{14}\\ =\frac{10d^2}{4c}\times\frac{14}{cd^3}\\ =\frac{2\times5\times d^2\times 2\times 7}{4\times c\times c\times d^2}\\ =\]

OpenStudy (bananas1234):

35/c^2d?

OpenStudy (unklerhaukus):

yes!, and you even ignored my typo,

OpenStudy (bananas1234):

great! :D thank you!